### v2

@@ 3.3.2 From WCP to Tutte Polynomial
Let $S$ be an infinite set of undirected graphs. Suppose that there is a fixed $\text{C}^{2}$ sentence $\Psi$ with a distinguished symmetric binary relation $E$ such that for each graph $G\in S$ of size $n$, there is a $\text{C}^{2}$ sentence $\Psi_{G}$ obtained by conjuncting $\Psi$ with cardinality constraints and ground unary literals, and $G$ is isomorphic to $G(\mu_{E})$ for each model $\mu\in\mathcal{M}_{\Psi_{G},n}$. [^10] For instance, [KSS22] gave the characterization of such graphs when the sentence is restricted to be in $\text{C}^{2}$. Note that in their work the counting parameters and the size of the sentence are unbounded while in our case they should be constant to apply the fast calculation of WCP. Then the Tutte polynomial of any $G\in S$ of size $n$ can be computed as follows in time polynomial in $n$:
$T(x,y)=\frac{f_{n}\left((x-1)(y-1)-1,\sqrt{y-1};\Psi_{T}(G),\mathds{1},\mathds{1},R\right)}{(x-1)^{cc(\mu_{E})}(y-1)^{n}\mathsf{WFOMC}(\Psi_{G},n,\mathds{1},\mathds{1})},$
@@
$f_{n}\left((x-1)(y-1)-1,\sqrt{y-1};\Psi_{T}(G),\mathds{1},\mathds{1},R\right)$
@@
By Proposition 3.5, computing $f_{n}\left((x-1)(y-1)-1,\sqrt{y-1};\Psi_{T}(G),\mathds{1},\mathds{1},R\right)$ is in polynomial time in $n$. Since the cardinality of $S$ is infinite, the size of $\Psi$ is independent of the size of $G$, thus $\mathsf{WFOMC}(\Psi_{G},n,\mathds{1},\mathds{1})$ is domain-liftable by the algorithm described in Section 2.4 as well. Therefore, we can obtain $T(x,y)$ in time polynomial in $n$. ∎
@@
$\frac{f_{n}\left((x-1)(y-1)-1,\sqrt{y-1};\Psi_{T}(G),\mathds{1},\mathds{1},R\right)}{(x-1)(y-1)^{n}}$
@@
$\Psi_{G}=$ $\Psi\land\bigwedge_{e=1}^{n}\left(Block_{v(e)}(e)\land\bigwedge_{j\neq v(e)}\lnot Block_{j}(e)\right)$
@@
$\frac{f_{n}\left((x-1)(y-1)-1,\sqrt{y-1};\Psi_{T}(G),\mathds{1},\mathds{1},R\right)}{(x-1)^{cc(G)}(y-1)^{n}}.$
@@
The above result answers positively the question in [vBK23] whether calculating the Tutte polynomial of a block-structured graph is hard. However, we remark that this result cannot be extended to computing the Tutte polynomial or evaluating any point for general graphs since representing a general graph by a first-order sentence involves a list of ground binary literals, making the WFOMC task #$\mathsf{\#P}$-hard [dBD12]. Therefore we are consistent with the hardness of computing the Tutte polynomial of general graphs [JVW90], while showing tractability for block-structured ones (with a constant number of blocks).
@@ 4.1 Definition
Let $\Psi$ be a first-order logic sentence possibly with cardinality constraints and ground unary literals, $w$ and $\overline{w}$ be two weighting functions and $R$ be a distinguished binary relation. The $n$-th Strict Strong Connectedness Polynomial (SSCP) $g_{n}(u,v)$ of $\Psi$ for the relation $R$ is the unique bivariate polynomial which satisfies:
@@
The $n$-th Non-strict Strong Connectedness Polynomial (NSCP) $\bar{g}_{n}(u,v)$ of $\Psi$ for the relation $R$ is the unique bivariate polynomial which satisfies:
@@
The predicates $A_{1},\cdots,A_{u}$ treat $R$ symmetrically and are defined similarly as those in WCP. As discussed in Proposition 3.2, we can imagine them as capturing the information about weakly connected components of $G(R)$. The predicates $B_{1},\cdots,B_{v}$ capture the directionality of edges in $G(R)$, enabling further directed axioms on $R$. We give an interpretation of both polynomials by the following proposition and present more properties and applications in Section 4.3.
@@
For the $n$-th SSCP and the $n$-th NSCP of the sentence $\Psi$, the weighting functions $w$ and $\overline{w}$ and the binary relation $R$, it holds that
@@
$g_{n}(u,v;\Psi,w,\overline{w},R)$ and $\bar{g}_{n}(u,v;\Psi,w,\overline{w},R)$ are bivariate polynomials in which both variables have a degree at most $n$.
@@
By Corollary 2.14 and 2.15, $\chi_{G(\mu_{R})}(v)$ and $\bar{\chi}_{G(\mu_{R})}(v)$ are polynomials of $v$ which has degree at most $n$. Since $cc(\mu_{R})\leq n$, the proposition follows immediately from Proposition 4.2. ∎
@@ 4.2 Fast Calculation
$C^{A}_{0}(x)$ =$=\lnot A_{1}(x)\wedge\lnot A_{2}(x)\wedge\dots\wedge\lnot A_{u}(x),$
$C^{B}_{0}(x)$ =$=\lnot B_{1}(x)\wedge\lnot B_{2}(x)\wedge\dots\wedge\lnot B_{v}(x),$
@@
$r_{(i,i^{\prime},i^{\prime\prime}),(j,j^{\prime},j^{\prime\prime})}=\begin{cases}r_{i,j}=\mathsf{WMC}(\psi_{i,j}(a,b),w,\overline{w}),&i^{\prime}=j^{\prime}\land i^{\prime\prime}=j^{\prime\prime},\\ r^{\to}_{i,j}=\mathsf{WMC}(\psi_{i,j}(a,b)\wedge\lnot R(b,a),w,\overline{w}),&i^{\prime}=j^{\prime}\land i^{\prime\prime}\neq j^{\prime\prime},\\ \begin{aligned} r^{\not-}_{i,j}=\mathsf{WMC}(&\psi_{i,j}(a,b)\wedge\lnot R(a,b)\\ &\wedge\lnot R(b,a),w,\overline{w}),\end{aligned}&i^{\prime}\neq j^{\prime},\\ \end{cases}$
@@
Figure 2: The 1-type structure of $\Psi_{R,u,v}$ which is a $(v+1)\times(u+1)$ grid. Each tiny gray box represents an original 1-type of $\Psi$. The bold gray cells (each containing $4\times 4$ boxes) refer to the replicas of the original 1-types. The DP algorithm proceeds first along the vertical direction (from cell $(4,0)$ to cell $(4,4)$) and then along the horizontal direction (from columns $0,\cdots,3$ to column $4$).
@@
Let us first consider the calculation along the vertical direction. Note that all columns in the grid are identical in terms of the weight of configurations, thus we only need to compute an arbitrary column $i^{\prime}\in[u]$. We define $h^{|}_{i^{\prime},\hat{v},\bm{c}}$ as the summed weight of 1-types $(\cdot,i^{\prime},0)\cdots,(\cdot,i^{\prime},\hat{v})$ where $\bm{c}=(c_{1},\cdots,c_{L})$ such that for each $i\in[L]$, $\sum_{i^{\prime\prime}=0}^{\hat{v}}c_{(i,i^{\prime},i^{\prime\prime})}=c_{i}$. Then we have
@@
$w^{|}_{in}(\bm{c^{*}})=\binom{|\bm{c}^{*}|}{\bm{c}^{*}}\prod_{i=1}^{L}W(C_{i},w,\overline{w})^{c^{*}_{i}}\cdot s_{i}^{\binom{c^{*}_{i}}{2}}\prod_{1\leq i<j\leq L}\left(r_{i,j}\right)^{c^{*}_{i}\cdot c^{*}_{j}}$ (5)
is the weight within cell $(i^{\prime},\hat{v})$, and
$w^{|}_{cross}(\bm{\bar{c}},\bm{c^{*}})=\prod_{i=1}^{L}\prod_{j=1}^{L}\left(r_{i,j}^{\to}\right)^{\bar{c}_{i}\cdot c^{*}_{j}}$
is the weight between cells $(i^{\prime},0),\cdots,(i^{\prime},\hat{v}-1)$ and cell $(i^{\prime},\hat{v})$. The initial value of $h^{|}_{i^{\prime},0,\bm{c}}$ is $w_{in}^{|}(\bm{c})$ for each $|\bm{c}|\leq n$.
In the same spirit, we can calculate the weight of the 1-type configurations in the horizontal direction. We define $h^{-}_{\hat{u},\bm{c}}$ as the summed weight of 1-types $(\cdot,0,\cdot),\cdots,(\cdot,\hat{u},\cdot)$ where $\bm{c}=(c_{1},\cdots,c_{L})$ such that for each $i\in[L]$, $\sum_{i^{\prime}=0}^{\hat{u}}\sum_{i^{\prime\prime}=0}^{v}c_{(i,i^{\prime},i^{\prime\prime})}=c_{i}$. By DP, the value of $h^{-}_{\hat{u},\bm{c}}$ can be calculated by
@@
$w^{-}_{cross}(\bm{\bar{c}},\bm{c^{*}})=\prod_{i=1}^{L}\prod_{j=1}^{L}\left(r_{i,j}^{\not-}\right)^{\bar{c}_{i}\cdot c^{*}_{j}}$
is the weight between columns $10,\cdots,\hat{u}-$ 1 and column $\hat{u}$. One can easily check that the initial value of $h^{-}_{0,\bm{c}}$ is $h^{|}_{0,v,\bm{c}}$ for each $|\bm{c}|\leq n$. Finally, the WFOMC of $\Psi_{R,u,v}$ can be obtained by summation over all possible configurations $\bm{c}$:
@@
The overall calculation is summarized in Algorithm 2, where $\dots$) takes $(n+1)^{2}$ points as input and returns a bivariate polynomial with degree $n$ of each variable. Since $h^{|}_{i^{\prime},v,\bm{c}}$ are the same for each $i^{\prime}$, we omit the subscript $i^{\prime}$ in the algorithm.
@@
6             $h^{|}_{v,\bm{c}}\leftarrow\sum_{\bm{\bar{c}}+\bm{c^{*}}=\bm{c}}h^{|}_{v-1,\bm{\bar{c}}}\cdot\binom{|\bm{c}|}{|\bm{c}^{*}|}\cdot w^{|}_{in}(\bm{c^{*}})\cdot w^{|}_{cross}(\bm{\bar{c}},\bm{c^{*}})$triangleright We obtain NSCP if $w^{|}_{in}(\bm{c^{*}})$ is given by Equation 5, or SSCP if it is given by Equation 6
@@
$w^{|}_{in}(\bm{c^{*}})=\binom{|\bm{c}^{*}|}{\bm{c}^{*}}\prod_{\begin{subarray}{c}i=1,\\ \lnot R(x,x)\in C_{i}\end{subarray}}^{L}W(C_{i},w,\overline{w})^{c^{*}_{i}}\cdot\left(s^{\not-}_{i}\right)^{\binom{c^{*}_{i}}{2}}\prod_{\begin{subarray}{c}1\leq i<j\leq L,\\ \lnot R(x,x)\in C_{i},\lnot R(x,x)\in C_{j}\end{subarray}}\left(r^{\not-}_{i,j}\right)^{c^{*}_{i}\cdot c^{*}_{j}}$ (6)
@@ 4.3 Applications
Recall from Proposition 4.2 that $\bar{g}_{n}(u,v;\Psi,w,\overline{w},R)$ equals to the sum of $W(\mu,w,\overline{w})\cdot(u+1)^{cc(\mu_{R})}\cdot\bar{\chi}_{G(\mu_{R})}(v+1)$ over all $\mu\in\mathcal{M}_{\Psi,n}$. The term $[u]\bar{g}_{n}(u-1,v;\Psi,w,\overline{w},R)$ only reserves the weight of models $\mu$ where $G(\mu_{R})$ is weakly connected.
@@
The proof directly follows from [MBS23, Proposition 14 and 15] that reduce WFOMC of $\text{C}^{2}$ with cardinality constraints and a single directed tree axiom or a single directed forest axiom to WFOMC of $\text{C}^{2}$ with cardinality constraints and a single acyclicity axiom. ∎
@@
Though some of the axioms have been proven to be domain-liftable such as the acyclicity axiom in [MBS23] and the linear order axiom in [TK23], by our approach we can further generalize these axioms to obtain more features such as restricting the number of weakly connected components of $G(R)$ of an acyclic graph. For example, if we require $G(R)$ to be the union of $k$ disjoint directed acyclic graphs, we can obtain the WFOMC by $(-1)^{n}\cdot[u^{k}]g_{n}(u-1,-2;\Psi,w,\overline{w},R)$.
@@ 5 Combinations of Axioms
An important advantage of the proposed approach over the prior works is that it can easily extend to new axioms by combining existing axioms. For instance, consider the polytree axiom that requires the interpretation of a distinguished binary predicate to be a polytree, i.e., a DAG whose underlying undirected graph is a tree. Adopting the methodology of existing research [vBK23, TK23, MBS23], one would typically delve into the literature on polytree enumeration and attempt to adapt the techniques (if they exist) to the WFOMC problem. However, even if such techniques are available, integrating them into current WFOMC algorithms is not a trivial task. In contrast, our approach leverages a single polynomial to encapsulate the structural characteristics of the models. When computing WFOMC with an axiom, the process involves extracting the necessary information from the polynomials, either by algebraically manipulating the coefficients or by evaluating the polynomial at specific points. Thus, to support a new axiom that is expressible as a combination of existing axioms, we can directly and synthetically apply the existing retrieval techniques. This streamlined approach eliminates the need for further algorithmic development for those axioms that can be captured by one of the polynomials.
@@
and let $f^{\prime}_{n}$ be the extended WCP defined in Definition 3.8. Then we have
@@
$\mathsf{WFOMC}(\Psi\land forest_{k}(R),n,w,\overline{w})$ $=[u^{k}v^{2(n-k)}]f^{\prime}_{n}(u-1,v;\Psi_{R},w,\overline{w},R),$

### v3

@@ 3.3.2 From WCP to Tutte Polynomial
Let $S$ be an infinite set of undirected graphs. Suppose that there is a fixed $\text{C}^{2}$ sentence $\Psi$ with a distinguished symmetric binary relation $E$ such that for each graph $G\in S$ of size $n$, there is a $\text{C}^{2}$ sentence $\Psi_{G}$ obtained by conjuncting $\Psi$ with cardinality constraints and ground unary literals, and $G$ is isomorphic to $G(\mu_{E})$ for each model $\mu\in\mathcal{M}_{\Psi_{G},n}$. [^10] For instance, [C2graphs] gave the characterization of such graphs when the sentence is restricted to be in $\text{C}^{2}$. Note that in their work the counting parameters and the size of the sentence are unbounded while in our case they should be constant to apply the fast calculation of WCP. Then the Tutte polynomial of any $G\in S$ of size $n$ can be computed as follows in time polynomial in $n$:
$T(x,y)=\frac{\hat{f}_{n}\left((x-1)(y-1)-1,\sqrt{y-1};\Psi_{T}(G),\mathds{1},\mathds{1},R\right)}{(x-1)^{cc(\mu_{E})}(y-1)^{n}\mathsf{WFOMC}(\Psi_{G},n,\mathds{1},\mathds{1})},$
@@
$\hat{f}_{n}\left((x-1)(y-1)-1,\sqrt{y-1};\Psi_{T}(G),\mathds{1},\mathds{1},R\right)$
@@
By Corollary 3.9, computing $f_{n}\left((x-1)(y-1)-1,\sqrt{y-1};\Psi_{T}(G),\mathds{1},\mathds{1},R\right)$ is in polynomial time in $n$. Since the cardinality of $S$ is infinite, the size of $\Psi$ is independent of the size of $G$, thus $\mathsf{WFOMC}(\Psi_{G},n,\mathds{1},\mathds{1})$ can be computed in time polynomial in $n$ by the algorithm described in Section 2.4 as well. Therefore, we can obtain $T(x,y)$ in time polynomial in $n$. ∎
@@
$\frac{\hat{f}_{n}\left((x-1)(y-1)-1,\sqrt{y-1};\Psi_{T}(G),\mathds{1},\mathds{1},R\right)}{(x-1)(y-1)^{n}}$
@@
$\Psi_{G}=$ $\Psi\land\bigwedge_{e\in V}\left(Block_{v(e)}(e)\land\bigwedge_{j\neq v(e)}\lnot Block_{j}(e)\right)$
@@
$\frac{\hat{f}_{n}\left((x-1)(y-1)-1,\sqrt{y-1};\Psi_{T}(G),\mathds{1},\mathds{1},R\right)}{(x-1)^{cc(G)}(y-1)^{n}}.$
@@
The above result answers positively the question in [WFOMC-tree-axioms] whether calculating the Tutte polynomial of a block-structured graph is hard. However, we remark that this result cannot be extended to computing the Tutte polynomial or evaluating any point for general graphs since representing a general graph by a first-order sentence involves a list of ground binary literals, making the WFOMC task $\mathsf{\#P}$-hard [WFOMC-conditioning]. Therefore we are consistent with the hardness of computing the Tutte polynomial of general graphs [tuttepoly-hardness], while showing tractability for block-structured ones (with a constant number of blocks).
@@ 4.1 Definition
Let $\Psi$ be a first-order logic sentence possibly with cardinality constraints and ground unary literals, $w$ and $\overline{w}$ be two weighting functions and $R$ be a distinguished binary relation. The $n$-th Strict Strong Connectedness Polynomial (SSCP) $g_{n}(u,v)$ of $\Psi$ for the relation $R$ is the bivariate polynomial which satisfies:
@@
The $n$-th Non-strict Strong Connectedness Polynomial (NSCP) $\bar{g}_{n}(u,v)$ of $\Psi$ for the relation $R$ is the bivariate polynomial which satisfies:
@@
The predicates $A_{1},\cdots,A_{u}$ treat $R$ symmetrically and are defined similarly as those in WCP. As discussed in Proposition 3.2, we can imagine them as capturing the information about weakly connected components of $G(R)$. The predicates $B_{1},\cdots,B_{v}$ capture the directionality of edges in $G(R)$, enabling further directed axioms on $R$. We give an interpretation of both SSCP and NSCP by the following proposition and present more properties and applications in Section 4.3.
@@
For the sentence $\Psi$, the weighting functions $w$ and $\overline{w}$ and the binary relation $R$, it holds that
@@
$g_{n}(u,v;\Psi,w,\overline{w},R)$ and $\bar{g}_{n}(u,v;\Psi,w,\overline{w},R)$ always exist, and they are unique bivariate polynomials in which both variables have a degree at most $n$.
@@
The existence of SSCP and NSCP follow immediately from Proposition 4.2. By Corollary 2.14 and 2.15, $\chi_{G(\mu_{R})}(v)$ and $\bar{\chi}_{G(\mu_{R})}(v)$ are polynomials of $v$ which has degree at most $n$. Since $cc(\mu_{R})\leq n$, the upper bounds of degree follow immediately from Proposition 4.2. ∎
@@ 4.2 Fast Calculation
$C^{A}_{0}(x)$ $=\lnot A_{1}(x)\wedge\lnot A_{2}(x)\wedge\dots\wedge\lnot A_{u}(x),$
$C^{B}_{0}(x)$ $=\lnot B_{1}(x)\wedge\lnot B_{2}(x)\wedge\dots\wedge\lnot B_{v}(x),$
@@
[…]
@@
Figure 2: The 1-type structure of $\Psi_{R,u,v}$ which is a $(v+1)\times(u+1)$ grid. Each tiny gray box represents an original 1-type of $\Psi$. The bold gray cells (each containing $4\times 4$ boxes) refer to the replicas of the original 1-types. The DP algorithm proceeds first along the vertical direction (from cell $(4,0)$ to cell $(4,4)$) and then along the horizontal direction (from columns $0,\cdots,3$ to column 44).
@@
[…]
@@
$w^{|}_{in}(\bm{c^{*}})=\binom{|\bm{c}^{*}|}{\bm{c}^{*}}\prod_{i=1}^{p}W(C_{i},w,\overline{w})^{c^{*}_{i}}\cdot s_{i}^{\binom{c^{*}_{i}}{2}}\prod_{1\leq i<j\leq p}\left(r_{i,j}\right)^{c^{*}_{i}\cdot c^{*}_{j}}$ (5)
[…]
$w^{|}_{cross}(\bm{\bar{c}},\bm{c^{*}})=\prod_{i=1}^{p}\prod_{j=1}^{p}\left(r_{i,j}^{\to}\right)^{\bar{c}_{i}\cdot c^{*}_{j}}$
[…]
@@
$w^{-}_{cross}(\bm{\bar{c}},\bm{c^{*}})=\prod_{i=1}^{p}\prod_{j=1}^{p}\left(r_{i,j}^{\not-}\right)^{\bar{c}_{i}\cdot c^{*}_{j}}$
is the weight between columns $0,\cdots,\hat{u}-1$ and column $\hat{u}$. One can easily check that the initial value of $h^{-}_{0,\bm{c}}$ is $h^{|}_{0,v,\bm{c}}$ for each $|\bm{c}|\leq n$. Finally, the WFOMC of $\Psi_{R,u,v}$ can be obtained by summation over all possible configurations $\bm{c}$:
@@
The overall calculation is summarized in Algorithm 2, where interpolate_2d($\dots$) takes $(n+1)^{2}$ points as input and returns a bivariate polynomial with degree $n$ of each variable. Since $h^{|}_{i^{\prime},v,\bm{c}}$ are the same for each $i^{\prime}$, we omit the subscript $i^{\prime}$ in the algorithm.
@@
6     $h^{|}_{v,\bm{c}}\leftarrow\sum_{\bm{\bar{c}}+\bm{c^{*}}=\bm{c}}h^{|}_{v-1,\bm{\bar{c}}}\cdot\binom{|\bm{c}|}{|\bm{c}^{*}|}\cdot w^{|}_{in}(\bm{c^{*}})\cdot w^{|}_{cross}(\bm{\bar{c}},\bm{c^{*}})$ $\triangleright$ We obtain NSCP if $w^{|}_{in}(\bm{c^{*}})$ is given by Equation 5, or SSCP if it is given by Equation 6
@@
$w^{|}_{in}(\bm{c^{*}})=\binom{|\bm{c}^{*}|}{\bm{c}^{*}}\prod_{\begin{subarray}{c}i=1,\\ \lnot R(x,x)\in C_{i}\end{subarray}}^{p}W(C_{i},w,\overline{w})^{c^{*}_{i}}\cdot\left(s^{\not-}_{i}\right)^{\binom{c^{*}_{i}}{2}}\prod_{\begin{subarray}{c}1\leq i<j\leq p,\\ \lnot R(x,x)\in C_{i},\lnot R(x,x)\in C_{j}\end{subarray}}\left(r^{\not-}_{i,j}\right)^{c^{*}_{i}\cdot c^{*}_{j}}$ (6)
@@ 4.3 Applications
Recall from Proposition 4.2 that $\bar{g}_{n}(u,v;\Psi,w,\overline{w},R)$ equals to the sum of $W(\mu,w,\overline{w})\cdot(u+1)^{cc(\mu_{R})}\cdot\bar{\chi}_{G(\mu_{R})}(v+1)$ over all $\mu\in\mathcal{M}_{\Psi,n}$. The term $u]\bar{g}_{n}(u-1,v;\Psi,w,\overline{w},R)$ only reserves the weight of models $\mu$ where $G(\mu_{R})$ is weakly connected.
@@
The proof directly follows from [WFOMC-axioms, Proposition 14 and 15] that reduce WFOMC of $\text{C}^{2}$ with cardinality constraints and a single directed tree axiom or a single directed forest axiom to WFOMC of $\text{C}^{2}$ with cardinality constraints and a single acyclicity axiom. ∎
@@
Though some of the axioms have been proven to be domain-liftable such as the acyclicity axiom in [WFOMC-axioms] and the linear order axiom in [WFOMC-linearorder-axiom], by our approach we can further generalize these axioms to obtain more features such as restricting the number of weakly connected components of $G(R)$ of an acyclic graph. For example, if we require $G(R)$ to be the union of $k$ disjoint directed acyclic graphs, we can obtain the WFOMC by $(-1)^{n}\cdot[u^{k}]g_{n}(u-1,-2;\Psi,w,\overline{w},R)$.
@@ 5 Combinations of Axioms
An important advantage of the proposed approach over the prior works is that it can easily extend to new axioms by combining existing axioms. For instance, consider the polytree axiom that requires the interpretation of a distinguished binary predicate to be a polytree, i.e., a DAG whose underlying undirected graph is a tree. Adopting the methodology of existing research [WFOMC-tree-axioms, WFOMC-linearorder-axiom, WFOMC-axioms], one would typically delve into the literature on polytree enumeration [polytree] and attempt to adapt the techniques (if they exist) to the WFOMC problem. However, even if such techniques are available, integrating them into current WFOMC algorithms is not a trivial task. In contrast, our approach leverages a single polynomial to encapsulate the structural characteristics of the models. When computing WFOMC with an axiom, the process involves extracting the necessary information from the polynomials, either by algebraically manipulating the coefficients or by evaluating the polynomial at specific points. Thus, to support a new axiom that is expressible as a combination of existing axioms, we can directly and synthetically apply the existing retrieval techniques. This streamlined approach eliminates the need for further algorithmic development for those axioms that can be captured by one of the polynomials.
@@
and let $\hat{f}_{n}$ be the extended WCP defined in Definition 3.8. Then we have
@@
$\mathsf{WFOMC}(\Psi\land forest_{k}(R),n,w,\overline{w})$ $=[u^{k}v^{2(n-k)}]\hat{f}_{n}(u-1,v;\Psi_{R},w,\overline{w},R),$